OpenStudy (anonymous):

How do you find the x-intercepts of x^(4)-8x^(2)+8?

OpenStudy (whpalmer4):

\[x^4-8x^2+8=0\] Why don't you try substitution of \[u = x^2\] to turn it into a quadratic: \[u^2-8u+8=0\]Solve that by your favorite method, then undo the substitution.

OpenStudy (anonymous):

I got it, thank you!

OpenStudy (whpalmer4):

What did you get for your answers?

OpenStudy (anonymous):

\[x=-\sqrt{4-\sqrt{2}}, \sqrt{{4}-\sqrt{2}}, -\sqrt{{4}+\sqrt{2}},\sqrt{{4}+\sqrt{2}}\]

OpenStudy (whpalmer4):

Oops, not so fast...those aren't quite correct.

OpenStudy (anonymous):

What did I do wrong?

OpenStudy (whpalmer4):

Should be a 2 in front of all of the \(\sqrt{2}\)'s there

OpenStudy (whpalmer4):

\[x = \pm\sqrt{4\pm2\sqrt{2}}\]

OpenStudy (anonymous):

Oh yeah, I just forgot to write in the 2 before the square root of 2. I have it in my work though.

OpenStudy (whpalmer4):

Ah, okay. We'll let you keep that medal ;-) I was worried that you might think there were only 2 solutions, so when I saw you had 4, I gave you the medal before I checked them closely :-)

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (whpalmer4):

you're welcome!